the speed of a stream is 5 mph. If a boat travels 66 miles downstream in the same time that it takes to travel 33 miles upsteam, what is the speed of the boat in still water?
Answer by mananth(16946)

Let the speed of boat in still water be x mph
upstream it will be x-5
down stream it will be x+5 mph
Time taken is ame
33/x= 66 / x+5
33(x+5)=66x
33x+165=66x
30x= 165
x=165/30
5.5 mph is the boat speed in still water